OpenStudy (jkbo):

Please help Which determinant has a value of 30? a.det [3 1] [3 9] b.det [3 1] [-3 9] c.det [3 -1] [-3 9] d.det [3 -1] [3 -9]

OpenStudy (bahrom7893):

Do you know how to find the determinant?

OpenStudy (jkbo):

Yes 3x3-9x1 but its always 0

OpenStudy (bahrom7893):

3*9-3*1 = 27-3 = 24 Not always 0

OpenStudy (bahrom7893):

What's 3*9 - (-3)*1?

OpenStudy (jkbo):

30

OpenStudy (anonymous):

b.det [3 1] [-3 9]

OpenStudy (jkbo):

thank you both